# Towards Unifying Understanding and Generation in the Era of Vision Foundation Models: A Survey from the Autoregression Perspective

### v1

@@ 2 Preliminary
PUMA [46] LQ Casual arXiv 2024
Janus [47] DP+VQ Casual arXiv 2024
SAR [48] VQ Casual+Bidirectional arXiv 2024
HART [49] VQ+LQ Casual arXiv 2024
@@
Baichuan-Omni [51] DP Casual arXiv 2024
ARIA [52] DP Casual arXiv 2024
Qwen2-VL [53] DP Casual arXiv 2024
Loong [54] VQ Casual arXiv 2024
Emu3 [55] VQ Casual arXiv 2024
MIO [56] LQ Casual arXiv 2024
G3PT [57] VQ Casual arXiv 2024
Open-MAGVIT2 [58] VQ Casual arXiv 2024
VILA-U [59] VQ Casual arXiv 2024
xGen-MM [60] DP Casual arXiv 2024
Lumina-mGPT [61] VQ Casual arXiv 2024
MARs [62] VQ Casual arXiv 2024
OmniTokenizer [63] VQ+LQ Casual arXiv 2024
GenLLaVA [64] LQ Casual arXiv 2024
EVE [65] DP Casual arXiv 2024
iVideoGPT [66] VQ Casual arXiv 2024
InternVL2 [67] LQ Casual CVPR 2024
VisionLLM v2 [68] LQ Casual NeurIPS 2024
@@
VAR [70] VQ Casual NeurIPS 2024
LlamaGen [71] VQ Casual arXiv 2024
Anole [72] VQ Casual arXiv 2024
Chameleon [73] VQ Casual arXiv 2024
Libra [74] VQ Casual ICML 2024
X-VILA [75] LQ Casual arXiv 2024
WorldGPT [76] DP Casual arXiv 2024
SEED-X [77] LQ Casual arXiv 2024
GiT [78] VQ Casual ECCV 2024
AnyGPT [79] LQ Casual arXiv 2024
DeLVM [80] VQ Casual ICML 2024
VideoLaVIT [81] VQ Casual ICML 2024
@@
Unified-IO 2 [82] VQ Casual CVPR 2024
Emu2 [83] LQ Casual CVPR 2024
VL-GPT [84] LQ Casual arXiv 2023
LVM [24] VQ Casual CVPR 2024
CoDi-2 [85] LQ Casual CVPR 2024
Mirasol3B [86] DP Casual CVPR 2024
TEAL [87] VQ Casual arXiv 2023
SEED-LLaMA [88] LQ Casual ICLR 2024
InstructSeq [89] VQ Casual arXiv 2023
LaViT [90] VQ Casual ICLR 2024
DreamLLM [19] LQ Casual ICLR 2024
Emu [91] LQ Casual ICLR 2024
CM3Leon [92] VQ Casual NeurIPS 2023
GILL [93] LQ Casual NeurIPS 2023
Parti [94] VQ Casual TMLR 2022
DALL-E [95] VQ Casual ICML 2021
Table 1: The overview of autoregressive vision foundation models. In the category of vision tokenizer, VQ denotes vector quantization, DP denotes direct projection, LQ denotes learnable query. In the category of autoregression backbone, Casual denotes casual Transformer ,Bidirectional denotes bidirectional Transformer, Prefix denotes prefix Transformer. + denotes both approaches are related.
@@ 3.2.1 Autoregression Transformer
Casual Transformer.
As shown in Fig. 2 (a), the casual attention utilizes masks to ensure that each token can only attend to itself and the previous tokens in the sequence [46, 47, 49, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 24, 85, 86, 87, 88, 89, 90, 19, 91, 92, 93, 94, 95]. This restriction aligns intuitively with autoregression. Formally, the representation of the $i$-th token $y_{i}$ in a self-attention layer is computed as:
@@
where $a_{ik}$ is the attention weight, indicating how much attention the $i$-th token gives to the $k$-th token. $v_{k}$ is the value embedding for the $k$-th token. In casual Transformer, masks are applied to attention weights, where $a_{ik}=0$ for $k>i$ and $\sum_{k=1}^{K}a_{ik}=1$. During training, the input sequence is prepended with a start token $[s]$, and the model attempts to predict the corresponding sequence shifted one token to the left. The loss is computed for all tokens in the sequence:
@@
As shown in Fig. 2 (b), the MAE-style autoregressive variant, named as MAR [69], is proved that bidirectional attention can also perform autoregression [45, 50]. Specifically, MAR reorders all tokens, and places unmasked tokens before the masked tokens. In this way, the masked tokens, i.e., the tokens to be predicted, are incidentally constrained to only access the previous tokens, as they are positioned at the end. Therefore, bidirectional attention allows each token to attend to all other tokens, and there is no need for casual masks to block attention weights. $a_{ik}$ in Eq. (8) satisfies $\sum_{k=1}^{K}a_{ik}=1$. During training, masked tokens $[m]$ are added at the end in the middle layer, and are the only ones in the sequence for which the loss needs to be computed as:
@@
where $M$ is the index set of masked tokens. Note that our analysis is based on the next token prediction. Models like MAR etc.can predict multiple masked tokens in parallel, but masked tokens are not visible to each other, regardless of whether they have already been predicted. They can only attend to previous unmasked tokens. SAR [48] further removes this limitation by allowing unknown tokens, i.e., tokens waiting for prediction, to be dynamically divided into multiple sets and performing next set prediction, enabling unpredicted sets to observe predicted sets and the initially known tokens. They achieve this through a Fully Masked Transformer, which can be viewed as a hybrid approach combining casual and bidirectional Transformer.
@@
As shown in Fig. 2 (c), some models adopt prefix attention [41], which also combines casual attention with bidirectional attention. In prefix Transformer, a prefix with the length of $S$ is treated with bidirectional attention, while the remainder of the sequence is processed with casual attention. For tokens in the prefix, i.e., $i\leq S$, $a_{ik}$ in Eq. (8) satisfies $a_{ik}=0$ for $k>S$ and $\sum_{k=1}^{S}a_{ik}=1$. For tokens outside the prefix, i.e., $i>S$, $a_{ik}$ in Eq. (8) satisfies $a_{ik}=0$ for $k>i$ and $\sum_{k=1}^{K}a_{ik}=1$. During training, only tokens beyond the prefix are computed for loss:
@@ 3.2.2 Autoregression Pattern
Figure 2: The overview of Autoregression Architecture. (a) Casual attention in casual Transformer. (b) Bidirectional attention in bidirectional Transformer. (c) Prefix attention in prefix Transformer. The token with $\downarrow$ is included in the loss calculation.

### v2

@@ 2 Preliminary
PUMA [46] LQ Causal arXiv 2024
Janus [47] DP+VQ Causal arXiv 2024
SAR [48] VQ Causal+Bidirectional arXiv 2024
HART [49] VQ+LQ Causal arXiv 2024
@@
Baichuan-Omni [51] DP Causal arXiv 2024
ARIA [52] DP Causal arXiv 2024
Qwen2-VL [53] DP Causal arXiv 2024
Loong [54] VQ Causal arXiv 2024
Emu3 [55] VQ Causal arXiv 2024
MIO [56] LQ Causal arXiv 2024
G3PT [57] VQ Causal arXiv 2024
Open-MAGVIT2 [58] VQ Causal arXiv 2024
VILA-U [59] VQ Causal arXiv 2024
xGen-MM [60] DP Causal arXiv 2024
Lumina-mGPT [61] VQ Causal arXiv 2024
MARs [62] VQ Causal arXiv 2024
OmniTokenizer [63] VQ+LQ Causal arXiv 2024
GenLLaVA [64] LQ Causal arXiv 2024
EVE [65] DP Causal arXiv 2024
iVideoGPT [66] VQ Causal arXiv 2024
InternVL2 [67] LQ Causal CVPR 2024
VisionLLM v2 [68] LQ Causal NeurIPS 2024
@@
VAR [70] VQ Causal NeurIPS 2024
LlamaGen [71] VQ Causal arXiv 2024
Anole [72] VQ Causal arXiv 2024
Chameleon [73] VQ Causal arXiv 2024
Libra [74] VQ Causal ICML 2024
X-VILA [75] LQ Causal arXiv 2024
WorldGPT [76] DP Causal arXiv 2024
SEED-X [77] LQ Causal arXiv 2024
GiT [78] VQ Causal ECCV 2024
AnyGPT [79] LQ Causal arXiv 2024
DeLVM [80] VQ Causal ICML 2024
VideoLaVIT [81] VQ Causal ICML 2024
@@
Unified-IO 2 [82] VQ Causal CVPR 2024
Emu2 [83] LQ Causal CVPR 2024
VL-GPT [84] LQ Causal arXiv 2023
LVM [24] VQ Causal CVPR 2024
CoDi-2 [85] LQ Causal CVPR 2024
Mirasol3B [86] DP Causal CVPR 2024
TEAL [87] VQ Causal arXiv 2023
SEED-LLaMA [88] LQ Causal ICLR 2024
InstructSeq [89] VQ Causal arXiv 2023
LaViT [90] VQ Causal ICLR 2024
DreamLLM [19] LQ Causal ICLR 2024
Emu [91] LQ Causal ICLR 2024
CM3Leon [92] VQ Causal NeurIPS 2023
GILL [93] LQ Causal NeurIPS 2023
Parti [94] VQ Causal TMLR 2022
DALL-E [95] VQ Causal ICML 2021
Table 1: The overview of autoregressive vision foundation models. In the category of vision tokenizer, VQ denotes vector quantization, DP denotes direct projection, LQ denotes learnable query. In the category of autoregression backbone, Causal denotes causal Transformer ,Bidirectional denotes bidirectional Transformer, Prefix denotes prefix Transformer. + denotes both approaches are related.
@@ 3.2.1 Autoregression Transformer
Causal Transformer.
As shown in Fig. 2 (a), the causal attention utilizes masks to ensure that each token can only attend to itself and the previous tokens in the sequence [46, 47, 49, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 24, 85, 86, 87, 88, 89, 90, 19, 91, 92, 93, 94, 95]. This restriction aligns intuitively with autoregression. Formally, the representation of the $i$-th token $y_{i}$ in a self-attention layer is computed as:
@@
where $a_{ik}$ is the attention weight, indicating how much attention the $i$-th token gives to the $k$-th token. $v_{k}$ is the value embedding for the $k$-th token. In causal Transformer, masks are applied to attention weights, where $a_{ik}=0$ for $k>i$ and $\sum_{k=1}^{K}a_{ik}=1$. During training, the input sequence is prepended with a start token $[s]$, and the model attempts to predict the corresponding sequence shifted one token to the left. The loss is computed for all tokens in the sequence:
@@
As shown in Fig. 2 (b), the MAE-style autoregressive variant, named as MAR [69], is proved that bidirectional attention can also perform autoregression [45, 50]. Specifically, MAR reorders all tokens, and places unmasked tokens before the masked tokens. In this way, the masked tokens, i.e., the tokens to be predicted, are incidentally constrained to only access the previous tokens, as they are positioned at the end. Therefore, bidirectional attention allows each token to attend to all other tokens, and there is no need for causal masks to block attention weights. $a_{ik}$ in Eq. (8) satisfies $\sum_{k=1}^{K}a_{ik}=1$. During training, masked tokens $[m]$ are added at the end in the middle layer, and are the only ones in the sequence for which the loss needs to be computed as:
@@
where $M$ is the index set of masked tokens. Note that our analysis is based on the next token prediction. Models like MAR etc.can predict multiple masked tokens in parallel, but masked tokens are not visible to each other, regardless of whether they have already been predicted. They can only attend to previous unmasked tokens. SAR [48] further removes this limitation by allowing unknown tokens, i.e., tokens waiting for prediction, to be dynamically divided into multiple sets and performing next set prediction, enabling unpredicted sets to observe predicted sets and the initially known tokens. They achieve this through a Fully Masked Transformer, which can be viewed as a hybrid approach combining causal and bidirectional Transformer.
@@
As shown in Fig. 2 (c), some models adopt prefix attention [41], which also combines causal attention with bidirectional attention. In prefix Transformer, a prefix with the length of $S$ is treated with bidirectional attention, while the remainder of the sequence is processed with causal attention. For tokens in the prefix, i.e., $i\leq S$, $a_{ik}$ in Eq. (8) satisfies $a_{ik}=0$ for $k>S$ and $\sum_{k=1}^{S}a_{ik}=1$. For tokens outside the prefix, i.e., $i>S$, $a_{ik}$ in Eq. (8) satisfies $a_{ik}=0$ for $k>i$ and $\sum_{k=1}^{K}a_{ik}=1$. During training, only tokens beyond the prefix are computed for loss:
@@ 3.2.2 Autoregression Pattern
Figure 2: The overview of Autoregression Architecture. (a) Causal attention in causal Transformer. (b) Bidirectional attention in bidirectional Transformer. (c) Prefix attention in prefix Transformer. The token with $\downarrow$ is included in the loss calculation.